40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4721     Entry time: Sun May 15 19:10:12 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Calibration of actuators : BS, ITMX and ITMY 

The AC response of the actuators on BS, ITMX and ITMY were re-measured by another technique.

Last time I estimated them by measuring the open-loop transfer functions, but this time the responses were measured in a more direct way.

The measured AC responses (60 Hz - 200 Hz) are :

      BS   = 1.643e-98 / f2  [m/counts] (corrected based on the plot below - Manasa)

     ITMX = 3.568e-9 / f2 [m/counts]

     ITMY = 3.542e-9 / f2 [m/counts]

Next : measurement of the PRM actuator response


(The technique) 

 This time a technique that Rana told me a week ago was used.

This technique allows us to directly measure the response of an actuator at high frequency without any loop corrections.

First of all, MICH has to be locked to keep MICH within the linear range of the error signal. So now MICH is a linear sensor to the mirror motions.

In the MICH control a steep low pass filter should be inserted in order to avoid unwanted effects from the control loop at the high frequencies.

For example I put a low pass filter composed of an elliptical filter whose cut-off frequency is at 50 Hz such that the control loop doesn't push the mirrors above the cut-off frequency.

Hence the error signal of MICH above 50 Hz directly corresponds to the motion of the mirrors including BS, ITMX and ITMY.

Taking a transfer function from an actuator to the MICH error signal directly gives the actuator response.

In my measurements MICH was locked by feeding the signal back to BS. The plot below is the expected open-loop transfer function for the MICH control.

oltf.png

You can see that the open loop TF suddenly drops above 50 Hz. The UGF was at about 20 Hz, confirmed by looking at the loop oscillation on DTT.

 

(Measurement)

 In the technique the error signal has to be calibrated to [m]. This time AS55_Q was used and calibrated based on a peak-to-peak measurement.

The peak to peak value in the MICH error signal was 8 counts, which corresponds to the sensor efficiency of 4.72e+07 [counts/m].

Then I took transfer functions from each suspension (i.e. C1:SUS-XXX_LSC_EXC) to the error signal at AS55_Q over a frequency range from 60 Hz to 200Hz.

For the transfer function measurements I ran the swept sine on DTT to get the data. Note that the PD whitening filters were on.

The plot below is the results of the measurements together with the fitting lines.

calib_actuators.png

In the fitting I excluded the data pints at 60 Hz, because their coherence was low due to the power line noise.

ELOG V3.1.3-